OpenStudy (anonymous):

Prove: cos 0 - cos 0•sin^2 0 = cos^3 0

OpenStudy (jdoe0001):

\(\bf cos(\theta)-cos(\theta)sin^2(\theta)=cos^3(\theta)?\)

OpenStudy (xapproachesinfinity):

just factor cos theta and the use the identity cos^2+sin^2=1